\begin{thebibliography}{99}


\bibitem {BJLPS} S.~Bates, W.~Johnson, J.~Lindenstrauss, D.~Preiss,
and G.~Schectman, {\it Affine approximation of Lipschitz functions
and nonlinear quotients}, Geometric and Functional Analysis {\bf 9}
(1999), 1092--1127.

\bibitem {B-Gr} R.~Beals and P.~Greiner, {\it Calculus on
Heisenberg Manifolds}, Annals of Mathematics Studies {\bf 119},
Princeton University Press, 1988.

\bibitem {BR} A.~Bella\"{\i}che and J.-J.~Risler, editors, {\it
Sub-Riemannian Geometry}, Birkh\"auser, 1996.

\bibitem {BL} Y.~Benyamini and J.~Lindenstrauss, {\it Geometric
Nonlinear Functional Analysis, Volume 1}, Colloquium Publications {\bf
48}, American Mathematical Society, 2000.

\bibitem {Bourdon} M.~Bourdon, {\it Immeubles hyperboliques,
dimension conforme et rigidit\'e de Mostow}, Geometric and Functional
Analysis {\bf 7} (1997), 245--268.

\bibitem {BP} M.~Bourdon and H.~Pajot, {\it Poincar\'e
inequalities and quasiconformal structure on the boundary of some
hyperbolic buildings}, Proceedings of the American Mathematical
Society {\bf 127} (1999), 2315--2324.

\bibitem {CW1} R.~Coifman and G.~Weiss, {\it Analyse Harmonique
Non-Commutative sur Certains Espaces Homog\`enes}, Lecture Notes in
Mathematics {\bf 242}, Springer-Verlag, 1971.

\bibitem {CW2} R.~Coifman and G.~Weiss, {\it Extensions of Hardy
spaces and their use in analysis}, Bulletin of the American
Mathematical Society {\bf 83} (1977), 569--645.

\bibitem {C-P} M.~Coornaert and A.~Papadopoulos, {\it Symbolic
Dynamics and Hyperbolic Groups}, Lecture Notes in Mathematics {\bf 1539},
Springer-Verlag, 1993.

\bibitem {D-S1} G.~David and S.~Semmes, {\it Fractured Fractals and
Broken Dreams: Self-Similar Geometry through Metric and Measure},
Oxford University Pres, 1997.

\bibitem {D-S2} G.~David and S.~Semmes, {\it Regular mappings between
dimensions}, Publicacions Matem\`atiques {\bf 44} (2000), 369--417.

\bibitem {DeVS} R.~DeVore and R.~Sharpley, {\it Maximal
Functions Measuring Smoothness}, Memoirs of the American Mathematical
Society {\bf 293}, 1984.

\bibitem {Duren} P.~Duren, {\it Theory of $H^p$ Spaces}, Academic
Press, 1970.

\bibitem {Falconer1} K.~Falconer, {\it The Geometry of Fractal Sets},
Cambridge University Press, 1985.

\bibitem {Falconer2} K.~Falconer, {\it Fractal Geometry: Mathematical
Foundations and Applications}, Wiley, 1990.

\bibitem {Falconer3} K.~Falconer, {\it Techniques in Fractal Geometry},
Wiley, 1997.

\bibitem {Federer} H.~Federer, {\it Geometric Measure Theory},
Springer-Verlag, 1969.

\bibitem {FS} G.~Folland and E.~Stein, {\it Hardy Spaces on
Homogeneous Groups}, Mathematical Notes {\bf 28}, Princeton University
Press, 1982.

\bibitem {F-S-SC} B.~Franchi, R.~Serapioni, and F.~Serra Cassano,
{\it Regular hypersurfaces, intrinsic perimeter and implicit function
theorem in Carnot groups}, to appear, Communications in Analysis and
Geometry.

\bibitem {misha-green} M.~Gromov, {\it Structures M\'etriques
pour les Vari\'et\'es Riemanniennes}, J.~Lafontaine and P.~Pansu,
editors, Cedic/Fernand Nathan, 1981.

\bibitem {misha-hypgroups} M.~Gromov, {\it Hyperbolic groups},
in {\it Essays in Group Theory}, S.~Gersten, editor, Mathematical
Sciences Research Institute Publications {\bf 8}, Springer-Verlag,
1987.

\bibitem {misha-asyminvgroups} M.~Gromov, {\it Asymptotic Invariants
of Infinite Groups}, Volume 2 of {\it Geometric Group Theory}, G.~Niblo
and A.~Roller, editors, Cambridge University Press, 1993.

\bibitem {misha-translation} M.~Gromov et al., {\it Metric
Structures for Riemannian and Non-Riemannian Spaces}, Birkh\"auser,
1999.

\bibitem {H-K} P.~Hajlasz and P.~Koskela, {\it Sobolev Met
Poincar\'e}, Memoirs of the American Mathematical Society {\bf 688},
2000.

\bibitem {H-W} W.~Hurewicz and H.~Wallman, {\it Dimension Theory},
Princeton University Press, 1941.

\bibitem {Hodges} W.~Hodges, {\it A Shorter Model Theory},
Cambridge University Press, 2002.

\bibitem {JLPS1} W.~Johnson, J.~Lindenstrauss, D.~Preiss, and
G.~Schechtman, {\it Uniform quotient mappings of the plane}, Michigan
Mathematical Journal {\bf 47} (2000), 15--31.

\bibitem {JLPS2} W.~Johnson, J.~Lindenstrauss, D.~Preiss, and
G.~Schechtman, {\it Almost Fr\'echet differentiability of Lipschitz
mappings between infinite-dimensional Banach spaces}, Proceedings of
the London Mathematical Society (3) {\bf 84} (2002), 711--746.

\bibitem {Kigami} J.~Kigami, {\it Analysis on Fractals},
Cambridge University Press, 2001.

\bibitem {Krantz1} S.~Krantz, {\it Function Theory of Several
Complex Variables}, Wiley, 1982.

\bibitem {Krantz2} S.~Krantz, {\it Lipschitz spaces, smoothness
of functions, and approximation theory}, Expositiones Mathematic{\ae}
{\bf 3} (1983), 193--260.

\bibitem {Krantz3} S.~Krantz, {\it Geometric Analysis and
Function Spaces}, Conference Board of the Mathematical Sciences
Regional Conference Series in Mathematics {\bf 81}, American
Mathematical Society, 1993.

\bibitem {K-P1} S.~Krantz and H.~Parks, {\it The Geometry of
Domains in Space}, Birkh\"auser, 1999.

\bibitem {K-P2} S.~Krantz and H.~Parks, {\it The Implicit Function
Theorem: History, Theory, and Applications}, Birkh\"auser, 2002.

\bibitem {Laakso1} T.~Laakso, {\it Ahlfors $Q$-regular spaces
with arbitrary $Q > 1$ admitting weak Poincar\'e inequality},
Geometric and Functional Analysis {\bf 10} (2000), 111--123; erratum,
{\bf 12} (2002), 650.

\bibitem {Laakso2} T.~Laakso, {\it Plane with $A_\infty$-weighted
metric not bi-Lipschitz embeddable to ${\bf R}^N$}, Bulletin of the
London Mathematical Society {\bf 34} (2002), 667--676.

\bibitem {MS} R.~Macias and C.~Segovia, {\it Lipschitz
functions on spaces of homogeneous type}, Advances in Mathematics {\bf
33} (1979), 257--270.

\bibitem {Valentino} V.~Magnani, {\it Elements of Geometric Measure
Theory on Sub-Riemannian Groups}, thesis, Scuola Normale Superiore,
Pisa, 2002.

\bibitem {Pertti} P.~Mattila, {\it Geometry of Sets and Measures in
Euclidean Spaces}, Cambridge University Press, 1995.

\bibitem {Mitchell} J.~Mitchell, {\it On Carnot--Carath\'eodory
metrics}, Journal of Differential Geometry {\bf 21} (1985), 35--45.

\bibitem {N-S} A.~Nagel and E.~Stein, {\it Lectures on Pseudodifferential
Operators: Regularity Theorems and Applications to Nonelliptic Problems},
Mathematical Notes {\bf 24}, Princeton University Press, 1979.

\bibitem {Ohshika} K.~Ohshika, {\it Discrete Groups},
Translations of Mathematical Monographs {\bf 207}, American
Mathematical Society, 2002.

\bibitem {Rudin1} W.~Rudin, {\it Principles of Mathematical
Analysis}, McGraw-Hill, 1976.

\bibitem {Rudin2} W.~Rudin, {\it Function Theory in the Unit
Ball of ${\bf C}^n$}, Springer-Verlag, 1980.

\bibitem {SS1} S.~Semmes, {\it Metric spaces and mappings
seen at many scales}, appendix in \cite{misha-translation}.

\bibitem {SS2} S.~Semmes, {\it Derivatives and difference quotients
for Lipschitz or Sobolev functions on various spaces}, in {\it Lecture
Notes on Analysis in Metric Spaces}, edited by L.~Ambrosio and
F.~Serra Cassano, 71--103, Scuola Normale Superiore, Pisa, 2000.

\bibitem {SS3} S.~Semmes, {\it Happy fractals and some aspects of
analysis on metric spaces}, submitted to Publicacions Matem\`atiques.

\bibitem {St1} E.~Stein, {\it Singular Integrals and
Differentiability Properties of Functions}, Princeton University
Press, 1970.

\bibitem {St2} E.~Stein, {\it Harmonic Analysis: Real-Variable
Methods, Orthogonality, and Oscillatory Integrals}, Princeton
University Press, 1993.

\bibitem {SW} E.~Stein and G.~Weiss, {\it Introduction to
Fourier Analysis on Euclidean Spaces}, Princeton University Press,
1971.

\bibitem {Strichartz1} R.~Strichartz, {\it Sub-Riemannian
Geometry}, Journal of Differential Geometry {\bf 24} (1986), 221--263;
corrections, {\bf 30} (1989), 595--596.

\bibitem {Strichartz2} R.~Strichartz, {\it Analysis on
fractals}, Notices of the American Mathematical Society {\bf 46}
(1999), 1199--1208.

\bibitem {TW} M.~Taibleson and G.~Weiss, {\it The molecular
characterization of certain Hardy spaces}, Ast\'erisque {\bf 77},
67--149, Soci\'et\'e Math\'ematique de France, 1980.

\bibitem {Taylor} M.~Taylor, {\it Noncommutative Microlocal Analysis I},
Memoirs of the American Mathematical Society {\bf 313}, 1984.

\bibitem {Zygmund} A.~Zygmund, {\it Trigonometric Series},
Volumes I and II, Cambridge University Press, 1979.



\end{thebibliography}

\documentclass[12pt,leqno,draft]{article}

\def\radius{\mathop{\rm radius}}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{sublemma}[theorem]{Sublemma}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{problem}[theorem]{Problem}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{claim}[theorem]{Claim}
\newtheorem{assumptions}[theorem]{Assumptions}
\newtheorem{examples}[theorem]{Examples}
\newtheorem{question}[theorem]{Question}
\newtheorem{sassumptions}[theorem]{Standing Assumptions}
\newtheorem{sassumption}[theorem]{Standing Assumption}
\newtheorem{conjecture}[theorem]{Conjecture}

\newcommand{\begintheorem}{\addtocounter{equation}{1}\begin{theorem}}
\newcommand{\beginlemma}{\addtocounter{equation}{1}\begin{lemma}}
\newcommand{\beginproposition}{\addtocounter{equation}{1}\begin{proposition}}
\newcommand{\beginsublemma}{\addtocounter{equation}{1}\begin{sublemma}}
\newcommand{\begindefinition}{\addtocounter{equation}{1}\begin{definition}}
\newcommand{\begincorollary}{\addtocounter{equation}{1}\begin{corollary}}
\newcommand{\beginproblem}{\addtocounter{equation}{1}\begin{problem}}
\newcommand{\beginremark}{\addtocounter{equation}{1}\begin{remark}}
\newcommand{\beginclaim}{\addtocounter{equation}{1}\begin{claim}}
\newcommand{\beginassumptions}{\addtocounter{equation}{1}\begin{assumptions}}
\newcommand{\beginexamples}{\addtocounter{equation}{1}\begin{examples}}
\newcommand{\beginquestion}{\addtocounter{equation}{1}\begin{question}}
\newcommand{\beginsassumptions}{\addtocounter{equation}{1}\begin{sassumptions}}
\newcommand{\beginsassumption}{\addtocounter{equation}{1}\begin{sassumption}}
\newcommand{\beginconjecture}{\addtocounter{equation}{1}\begin{conjecture}}

\begin{document}

\title{Some questions related to fractals}

\author{Stephen Semmes}

\date{}

\maketitle

\subsubsection*{Classes of fairly smooth functions}

	A fundamental aspect of harmonic analysis on Euclidean spaces
deals with various classes of real or complex-valued functions, such
as the $C^k$ classes of functions which are continuous and
continuously differentiable up to order $k$, where $k$ is a
nonnegative integer.  If $k = 0$, this simply means that the function
is continuous.  Of course continuity of a function makes sense on any
metric space, or on topological spaces more generally, while the
notion of \emph{derivatives} entails more structure.  Compare with
\cite{Rudin1}.

	We can be more precise and consider $C^{k, \alpha}$ classes of
functions, where $k$ is a nonnegative integer and $\alpha$ is a real
number, $0 \le \alpha \le 1$.  If $\alpha = 0$, then $C^{k, \alpha}$
is taken to mean the same as $C^k$; when $\alpha > 0$, $C^{k, \alpha}$
consists of the $C^k$ functions with the extra property that their
$k$th order derivatives are locally \emph{H\"older continuous of order
$\alpha$}.  For this let us recall that a function $h(x)$ on a metric
space $(M, d(x,y))$ is H\"older continuous of order $\alpha$ if there
is a nonnegative real number $A$ such that
\begin{equation}
\label{Holder continuity of order alpha}
	|h(x) - h(y)| \le A \, d(x,y)^\alpha \quad \hbox{for all } x, y \in M 
\end{equation}
This makes sense on any metric space, so that $C^{0, \alpha}$ classes
of functions make sense on any metric space, but $C^{k, \alpha}$
involves more structure when $k \ge 1$.

	The condition (\ref{Holder continuity of order alpha}) also
makes sense for $\alpha > 1$, but on a Euclidean space it would then
imply that the function is constant, because the first derivatives
would be identically equal to $0$.  One can also check this fact more
directly.  On a general metric space, this is not true in general.
However, it is true under fairly mild conditions, e.g., if the metric
space is connected and there are enough rectifiable curves around,
because such a function would have to be constant on each one of them.
Cantor sets and snowflake curves are basic examples of metric spaces
in which there are a lot of functions which satisfy (\ref{Holder
continuity of order alpha}) with $\alpha > 1$.

	Another smoothness class for functions on ${\bf R}^n$
is given by the \emph{Zygmund condition}
\begin{equation}
\label{Zygmund condition}
	|f(x + v) + f(x - v) - 2 f(x)| \le L \, |v| 
				\quad \hbox{for all } x, v \in {\bf R}^n,
\end{equation}
where $f$ is a continuous function on ${\bf R}^n$ and $L$ is a
nonnegative real number.  This condition is implied by (\ref{Holder
continuity of order alpha}) when $\alpha = 1$, with $L = 2 A$, but the
converse does not work, even locally.  It can be shown that if a
continuous function $f$ satisfies (\ref{Zygmund condition}), then it
is locally H\"older continuous of order $\alpha$ for each $\alpha <
1$.

	Note that the Zygmund condition is not defined on general
metric spaces.  Here is a nice reformulation of it, although not
exactly with the same constant, in terms of affine functions on ${\bf
R}^n$.  A function $f$ on ${\bf R}^n$ satisfies the Zygmund condition
for some $L$ if and only if there is a nonnegative real number $L'$
such that for every ball $B$ in ${\bf R}^n$ there is an affine
function $a(x)$ such that
\begin{equation}
\label{Zygmund condition, revisited}
	\sup_{x \in B} |f(x) - a(x)| \le L' \radius(B).
\end{equation}
It is very easy to go from this to (\ref{Zygmund condition}) with $L = L'$,
but the other direction is more complicated.

	If we replace $|v|$, $\radius(B)$ on the right sides of
(\ref{Zygmund condition}), (\ref{Zygmund condition, revisited}) with
$|v|^{1 + \alpha}$, $\radius(B)^{1 + \alpha}$, respectively, where $0
< \alpha \le 1$, then the corresponding conditions can be shown to be
equivalent to $f$ being continuously differentiable with first
derivatives H\"older continuous of order $\alpha$.  To go from
H\"older continuous first derivatives to the other conditions is
basically a matter of calculus, but the other direction is again more
complicated.  If one replaces $|v|$, $\radius(B)$ on the right sides
of (\ref{Zygmund condition}), (\ref{Zygmund condition, revisited})
with $|v|^{1 + \alpha}$, $\radius(B)^{1 + \alpha}$ when $\alpha > 1$,
then the resulting conditions imply that the second derivatives of $f$
are equal to $0$, and $f$ is affine.  To accommodate higher powers of
$|v|$ or $\radius(B)$, one can use differences of higher order in
place of the second difference in (\ref{Zygmund condition}), or
polynomials of higher degree in place of affine functions in
(\ref{Zygmund condition, revisited}).

	Some basic references concerning these matters are \cite{Duren,
Krantz2, MS, SS1, SS3, St1, SW, TW, Zygmund}.

	All of this uses the special structure of Euclidean spaces, to
define differences of order at least $2$, or to have affine functions
and polynomials of higher degree, rather than just constants.  What
about other metric spaces, and the structure that they might have?
This leads to a lot of questions.

	In the case of Heisenberg groups, other nilpotent Lie groups,
and sub-Riemannian spaces more generally, a lot of theory has been
developed along these lines.  There are various smoothness classes
analogous to classical ones on Euclidean spaces which are adapted to
the new geometry.  A basic notion is that $C^\infty$ functions or
polynomials in general might be the same as before, but more precise
degrees of smoothness are now different, and different kinds of
degrees of polynomials are used.  See \cite{FS, Krantz3}, for instance.

	What about fractal sets sitting inside of Euclidean spaces?
For these the usual affine functions and polynomials can be used,
just restricted to the fractal set under consideration.  One
can also define classes like $C^{k, \alpha}$ on the set by taking
restrictions of $C^{k, \alpha}$ functions from the ambient Euclidean
space.  Well-known extension theorems of Whitney characterize
functions on a closed set in a Euclidean space which can be extended
to $C^{k, \alpha}$ functions on the whole Euclidean space, as in
\cite{K-P1, St1}.

	Smoothness inherited from the ambient space in this manner
is quite different from the kind of regularity discussed in
\cite{Kigami, Strichartz2}, for instance.

	Other very interesting examples with their own special behavior
can be found in \cite{Bourdon, BP, Laakso1, Laakso2}.

\subsubsection*{Semi-Markovian spaces}

	Gromov \cite{misha-hypgroups, C-P} introduced a remarkable
notion of ``semi-Markovian spaces'' for which there are combinatorial
``presentations'', and which includes a number of standard fractals.
This appears to have a lot of room for interesting analysis of
functions.

\subsubsection*{Families of fractals}

	It seems to me that there are a lot of issues around the
general theme of ``families of fractals''.  A basic set-up would be to
have two metric spaces $T$ and $P$ and a mapping $\pi : T \to P$,
where the fibers $\pi^{-1}(p)$, $p \in P$, would be the fractals in
the family.  It is customary to call $T$ the ``total space'' for the
family of fractals, and $P$ the ``space of parameters''.  One might
ask that $\pi$ be Lipschitz, which is to say H\"older continuous of
oder $1$.  One could also ask for some nondegeneracy conditions, along
the lines of the nonlinear quotient mappings discussed in \cite{BJLPS,
BL, JLPS1, JLPS2}, or $(\tau, \rho)$-regular mappings or noncollapsing
mappings as in \cite{D-S2}.  For $(\tau, \rho)$-regular mappings,
$\tau$, $\rho$ could be some kind of dimensions for $T$, $P$,
respectively.

	Here is a somewhat more specific type of situation.  Suppose
that $n$ is a positive integer, and consider the Cartesian product
${\bf R}^n \times {\bf R}$.  Let $\lambda : {\bf R}^n \times {\bf R}
\to {\bf R}$ be the standard projection onto the last coordinate, so
that $\lambda(x, t) = t$ for all $x$ in ${\bf R}^n$ and $t$ in ${\bf
R}$.  Let $E$ be a subset of ${\bf R}^n \times {\bf R}$, and assume
for instance that $E$ is compact and that $\lambda(E) = [0,1]$.  The
sets $E(t) = \lambda^{-1}(t)$, $0 \le t \le 1$, would be the fractals
in the family.  Let $d$ be a positive real number, and suppose that
$E(t)$ has Hausdorff dimension $d$ for all $t$ in $[0,1]$, or almost
all $t$.  Then $E$ itself should have Hausdorff dimension at least $d
+ 1$, and if the Hausdorff dimension of $E$ is equal to $d + 1$,
then that reflects a kind of regularity in the situation.  Of course
there are a number of variants of this.

\end{document}